Prev Next

SW&IT Meeting Minutes 2021-12-06

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211129
LR20211129 - Bec, Matthieu 40.0
LR20211129 - Mechtley, Brandon 28.0
LR20211129 - Eychaner, Glenn 0.0
LR20211129 - Edmeades, David 40.0
LR20211129 - Grenz, Paul 16.0
LR20211129 - Hooper, Stephen 0.0
LR20211129 - Zhang, Yang 40.0
LR20211129 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • Upper Left TreeHouse PLC Network Connection (JPH)
  • microphone and speaker(s) to rm507-4 (or -1) (DT)
  • REBOOT obs4 (JCG)

Team reports

MB

  • Last week
    • OC, Management, ATT PM
    • Performance Appraisals
    • shark-nir PAE
    • WFS simulator, memleak check
  • This week
    • Performance Appraisals (end)
    • shark-nir PAE ctd.
    • WFS simulator ctd.
    • AO PB mini-crate, mixed-mode controls discussions with BM, GB ?

YZ

  • Previous:
    • Finished performance review with MB.
    • TMS meeting and some software update (e.g., the extraction utility).
    • MODS WCS program evaluation with Olga (overall we had pretty good understandings now, WCS database TTL limit temporarily extended).
    • New image utility program (analyzing and cataloging different LBT image types and some keywords).
    • LBC alerts script investigation (perhaps need refactor/cleaning).
    • Some MODS problems Jenny had encountered (moving things to obs5/6 gradually).
  • Next:
    • Finish MODS WCS program (and perhaps work on a centOS RPM build).
    • Continue new image utility work.
    • May tweak LBC alert scripts.
    • Checkout some MODS operations on obs5.
    • TMS meeting and software work.
    • Working on LBC bias adjustment wiki page.

DE

  • Last week:
    • Ordered laptop for JW
    • Reinstalled Win 10 on backup server with standard version that works with UITS' key.
    • Veeam back up and running, super easy to restore config
    • Pursued ticket with NetOps, they put a job in to activate AC's extra Ethernet ports
    • Syno drive swap, disk return
    • Another drive RMA, for some reason WD isn't allowing automated RMAs of some disks. I'll need to follow up with them.
    • xzhang-wfs clone for MB, should be done this morning
  • This week:
    • Tweak Veeam config to take advantage of new local storage
    • Second Syno drive swap
    • Talk to Ketiv about new Vault setup
    • archive-in iSCSI device
    • Vacation Thursday-Monday

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Created a pull request for the LBTO GitHub.
    • Completed a test version of the RTD with proper threading.
  • This week
    • Merge the pull request into the LBTO GitHub.
    • Create a release version (2.14.0) of the new LUCI software, and install it on luci.luci and in stow on the observer machines.
    • Test the new version of the RTD, and release the test version to SciOps.
    • Fix the error handling in the RTD so all errors go to the same place, rather than being spread across System.out, JOptionPanes, and the MessageServer GUI.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

BM

  • Last week
    • worked with JC on bin4 pupil files
    • keep_up fix--point to ~/live
    • PB upgrade meeting with GB/MB/XZ
    • New Housekeeper ready to test
    • Auxloops leaking multiprocessing temporary files--testing procedure is on Wiki.
    • LBTI offset interruption during SafeSkip issue--followed up with Amali/Alfio, may need to set up time to reproduce with some additional logging during offsets …
    • GH issue bookkeeping
  • This week
    • Keep looking into LBTI SafeSkip issue
    • Meet about "mixed mode" ASM operation with GB/MB
    • Questions for XZ about PB SW upgrade status, simulation mode
    • WfsArbitrator memory leak--MB investigating
    • GH issues:
      • ccd_viewer dying after StopAo
      • ResumeAo overwriting manually loaded gains
      • Python app logging, masterdiagnostic field width, etc.

PSG

  • Last week
    • Final testing of proxy.
    • Works correctly, allows ARGOS software and OSS to both talk to the HL2000.
    • Commands are not "mixed".
  • This week
    • Pull request to merge OSS code in with the next release?
    • Work on GCS.

SH

  • reboot web machine yesterday
  • VMWare (install partially)
  • Monitors for rm507, obs
  • Quotes for switches (due today)
  • broken monitors (talk to LeRoy Wednesday)

next:

  • 507 monitors (waiting on cables, network issues on newer 507-1).
  • VMware (finish install, e-mail Arto, etc).
  • ROFIN battery?
~
Topic revision: r9 - 06 Dec 2021, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback